Given: The area of the rectangle is 14cm2.
The width of a rectangle is 5 more than the length.
Find: The diagonal of the rectangle

Answer by addingup(3677)

length: x
width: x+5
x*x+5 = 14
x^2 = 9
x = 3 this is the length
x+5 = 3+5 = 8 is the width
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3^2+8^2 = diagonal^2
9+64 = 73 take the square root
sqrt 73 = 8.54 is the diagonal